OpenStudy (pokemon23):

A flyer on reconnaissance duty spent 5 hours on a mission . He flew out from his base with the wind at the rate of 360mph and returned to his base over the same route, flying against the wind at the rate of mph. How many miles did he fly out before he turned back?

OpenStudy (anonymous):

at the rate of mph? the second one?

OpenStudy (anonymous):

do you want to learn the material or just get a letter to type in to the computer?

OpenStudy (pokemon23):

i want to learn

OpenStudy (anonymous):

okay; do you know the basic rate equation, rate * time = distance?

OpenStudy (pokemon23):

please that will be good

OpenStudy (pokemon23):

yes

OpenStudy (pokemon23):

I know that rule

OpenStudy (anonymous):

ok. also i think you left one of the speeds (in mph) out of the problem?

OpenStudy (pokemon23):

let me check

OpenStudy (pokemon23):

A flyer on reconnaissance duty spent 5 hours on a mission . He flew out from his base with the wind at the rate of 360mph and returned to his base over the same route, flying against the wind at the rate of 240mph. How many miles did he fly out before he turned back?

OpenStudy (pokemon23):

My mistake*

OpenStudy (anonymous):

ok, so let's start writing some equations.

OpenStudy (pokemon23):

yes

OpenStudy (pokemon23):

so

OpenStudy (anonymous):

so first, let's call "t_out" the time he spent flying out, and "t_back" the time he spent flying back. we know: rate * time = distance, so on the way out, 360 * t_out = d similarly on the way back 240 * t_back = d The two d's are the same since he flew the same distance out and back. So we know: 360 * t_out = 240 * t_back. Does that make sense so far?

OpenStudy (pokemon23):

yes but t=time right

OpenStudy (anonymous):

well I split it into two times, the time he spent flying out and the time he spent coming back. the two times are not the same.

OpenStudy (pokemon23):

now we have to find t?

OpenStudy (anonymous):

there are two different t's in this problem. you use the rate equation separately on the outbound trip and the inbound trip.

OpenStudy (pokemon23):

360*t=360t 240*t=240t

OpenStudy (pokemon23):

we need to find how many miles did he fly out before he turned back?

OpenStudy (anonymous):

the two t's are different

OpenStudy (pokemon23):

ok.?

OpenStudy (anonymous):

t_out and t_back. they are not the same t.

OpenStudy (anonymous):

anyway you know that 360*t_out = 240*t_back, and t_out+tback = 5. now you simultaneously solve those two equations. do you know how to do that?

OpenStudy (pokemon23):

is this correct?

OpenStudy (anonymous):

no, that's not correct; that would be if he flew out for 5 hours and flew back for 5 hours. but that's not what happened. He flew for a *total* of 5 hours.

OpenStudy (anonymous):

You don't know how much of that 5 hours was spent going out, and how much coming back. that's what I'm trying to help you solve. 360 * t_out = 240*t_back t_out + t_back = 5 Can you simultaneously solve those two equations?

OpenStudy (pokemon23):

I'm not sure... but I think you subtract

OpenStudy (anonymous):

subtract what?

OpenStudy (pokemon23):

you said their both diffrent

OpenStudy (anonymous):

yes...? so what are you subtracting?

OpenStudy (pokemon23):

360-240?

OpenStudy (pokemon23):

I'm wrong?

OpenStudy (anonymous):

no. do you know how to solve systems of equations?

OpenStudy (pokemon23):

yes

OpenStudy (pokemon23):

But i need an equation to solve it?

OpenStudy (anonymous):

OK so you have to solve the following system of equations: 360 * t_out = 240*t_back t_out + t_back = 5 can you do that?

OpenStudy (pokemon23):

No ..

OpenStudy (anonymous):

ok so you don't know how to solve systems of equations?

OpenStudy (pokemon23):

No

OpenStudy (anonymous):

well, that's a big topic, but it's something you need to know how to do in order to solve this problem

OpenStudy (pokemon23):

I know you have to subtract a number to both sides to find the missing variable

OpenStudy (anonymous):

you haven't been doing linear equations in other parts of your class?

OpenStudy (anonymous):

Well if you simultaneously solve 360*t_out = 240*t_back t_out+tback = 5 t_out = 5-t_back 360*(5-t_back) = 240 * t_back t_back = 3 t_out = 2 d = t_out*360 = 720 there you go

OpenStudy (pokemon23):

just regular equations ;0

OpenStudy (anonymous):

A flyer on reconnaissance duty spent 5 hours on a mission . He flew out from his base with the wind at the rate of 360mph and returned to his base over the same route, flying against the wind at the rate of 240mph. How many miles did he fly out before he turned back? ------------------------------------------- r t = d Let t be the time to fly to the objective @ 360 mph. To fly back @ 240 mph requires (5-t) hours. The distance out is equal to the distance back. Solve the following for t: 360 t = 240 (5 - t) 360 t =1200 - 240 t 600 t = 1200 t = 2 360*2 = 720, the miles to fly out. 240(5-2) = 720. the miles to fly back.

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

whoa, i gave the answer like an hour ago and got no medal? I don't understand

OpenStudy (pokemon23):

sorry i was talking to my tutor hero

OpenStudy (pokemon23):

np

OpenStudy (pokemon23):

thanks for your answers

OpenStudy (pokemon23):

willing to help me in my other question/

OpenStudy (pokemon23):

yes???????

OpenStudy (anonymous):

One last comment.\[r*t = d, t=\frac{d}{r} \]One could solve for d directly using the fact that the sum of the times out and back is five hours.\[\frac{d}{360}+\frac{d}{240}=5 \]However, the arithemetic manipultion is a liitle bit more involved. For example one could multiply each side of the equation by the LCM[360,240]], the lowest common multiple, 720, to obtain after simplification: 5 d = 3600 d = 720 miles

OpenStudy (pokemon23):

lol thanks I thought you lag problems? at first in till I realized you had a response for me I appreciate your enthusiasm

OpenStudy (anonymous):

Thank you. I like to complete answers when possible.

OpenStudy (anonymous):

sorry. should have typed: "I like to provide complete answers when possible."

OpenStudy (pokemon23):

its alright can you go and help me in my other question

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Is is the triangle problem?

OpenStudy (pokemon23):

yes

OpenStudy (anonymous):

I'll go there now.

OpenStudy (pokemon23):

ty